# On cutoff via rigidity for high dimensional curved diffusions

### v3

@@ Abstract.
We consider overdamped Langevin diffusions in Euclidean space, with curvature equal to the spectral gap. This includes the Ornstein – Uhlenbeck process as well as non Gaussian and non product extensions with convex interaction, such as the Dyson process from random matrix theory. We show that a cutoff phenomenon or abrupt convergence to equilibrium occurs in high-dimension, at a critical time equal to the logarithm of the dimension divided by twice the spectral gap. This cutoff holds for Wasserstein distance, total variation, relative entropy, and Fisher information. A key observation is a relation to a spectral rigidity, linked to the presence of a Gaussian factor. A novelty is the extensive usage of functional inequalities, even for short-time regularization, and the reduction to Wasserstein. The proofs are short and conceptual. Since the product condition is satisfied, an Lp cutoff hold for all p. We moreover discuss a natural extension to Riemannian manifolds, a link with logarithmic gradient estimates in short-time for the heat kernel, and ask about stability by perturbation. Finally, beyond rigidity but still for diffusions, a discussion around the recent progresses on the product condition for nonnegatively curved diffusions leads us to introduce a new curvature product condition.
@@ 1.1. Diffusions on Euclidean spaces, with convex potential
1.2. Cutoff for high dimensional curved diffusions
@@
$\mathrm{W}_{2}(\nu,\mu)=\inf_{(X,Y)}\sqrt{\mathbb{E}(|X-Y|^{2})}$ (1.4)
where the infimum runs over all couples $(X,Y)$ of random variables with $X\sim\nu$ and $Y\sim\mu$, see Santambrogio (2015). By an abuse of notation, we also write $\mathrm{W}_{2}(Z,\mu)=\mathrm{W}_{2}(\mathrm{Law}(Z),\mu)$.
@@
$\frac{\mathrm{e}^{-\lambda_{1}t}}{\sqrt{\lambda_{1}}}\Bigr(\sup_{x_{0}\in S}\Lambda(x_{0})\Bigr)^{1/2}\leq\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t},\mu)\leq\mathrm{e}^{-\lambda_{1}t}\sup_{x_{0}\in S}\Bigr(\int{|x-x_{0}|^{2}\mathrm{d}\mu(x)}\Bigr)^{1/2}$ (1.6)
with, denoting $E_{1}$ the eigenspace of -$\mathcal{L}$ associated to $\lambda_{1}$, and $k_{1}=\dim(E_{1})$,
$\Lambda(x_{0})=k_{1}+\sup_{(f_{1},\ldots,f_{k_{1}})}\sum_{i=1}^{k_{1}}|f_{i}(x_{0})|^{2},$ (1.7)
where the supremum runs over the set of orthonormal bases of $E_{1}$. Moreover, if $\mu$ is centered and $S=\{x\in\mathbb{R}^{d}:|x|\leq R\}$, then $\sup_{x_{0}\in S}\Lambda(x)$ can be replaced by $k_{1}+\lambda_{1}R^{2}$.
Remark 1.2 (Dimensions).
We always have $k_{1}\leq d$, see (Barthe and Klartag, 2020, Lem. 14).
@@
$S^{(d)}=\{x\in\mathbb{R}^{d}:|x-m^{(d)}|\leq c\sqrt{d}\},\quad\text{for an arbitrary constant $c$}.$ (1.11)
@@
$\mathrm{H}(\nu\mid\mu)=\int\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\log\Bigr(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\Bigr)\mathrm{d}\mu=\int\log\Bigr(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\Bigr)\mathrm{d}\nu\in[0,+\infty]$ (1.13)
@@
$\mathrm{I}(\nu\mid\mu)=\int\frac{|\nabla\frac{\mathrm{d}\nu}{\mathrm{d}\mu}|^{2}}{\frac{\mathrm{d}\nu}{\mathrm{d}\mu}}\mathrm{d}\mu=\int\Bigr|\nabla\log\Bigr(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\Bigr)\Bigr|^{2}\mathrm{d}\nu\in[0,+\infty],$ (1.14)
with the convention $\mathrm{I}(\nu\mid\mu)=+\infty$ is $\nu$ is not absolutely continuous with respect to $\mu$. By an abuse of notation, we take the freedom of replacing $\nu$ by $X\sim\nu$ in these expressions.
@@
Under the setting of Corollary 1.3, for all $\varepsilon>0$,
@@
If $\pi$ and $\pi^{\perp}$ are the orthogonal projections on $\mathbb{R}(1,\ldots,1)$ and its orthogonal, respectively, then $|x|^{2}=|\pi(x)|^{2}+|\pi^{\perp}(x)|^{2}$, while the translation invariance of $W$ in the direction $(1/\sqrt{n},\ldots,1/\sqrt{n})$ gives $W(x)=W(\pi(x)+\pi^{\perp}(x))=W(\pi^{\perp}(x))$, therefore
@@
This covers as a special degenerate case the Dyson – OU or DOU process studied in Chafaï and Lehec (2020); Boursier et al. (2023), when
@@
the degeneracy being equivalent to define the Dyson – OU process on the convex domain $\{x\in\mathbb{R}^{d}:x_{1}>\cdots>x_{n}\}$ instead of on the whole space $\mathbb{R}^{d}$, to exploit convexity. In this case, the symmetric Hermite polynomial $x_{1}+\cdots+x_{d}$ is an eigenfunction associated to the spectral gap, the tip of an iceberg of integrability, as observed in Lassalle (1991).
@@ 1.3. Product condition and $L^{p}$ cutoff at the mixing time
$\lim_{d\to\infty}\sup_{x^{(d)}_{0}\in S^{(d)}}\Bigr\|\frac{\mathrm{d}\mu^{(d)}_{t}}{\mathrm{d}\mu^{(d)}}-1\Bigr\|_{\mathrm{L}^{p}(\mu^{(d)})}=\begin{cases}\max&\text{if $t_{d}=(1-\varepsilon)t_{*}$}\\ 0&\text{if $t_{d}=(1+\varepsilon)t_{*}$}\end{cases}$ (1.24)
where $\max=1$ if $p=1$ and $\max=+\infty$ if $p>1$, see Chafaï (2024) for a pedagogical presentation. We emphasize that Corollaries 1.3, 1.4, and 1.6 give the mixing time as well as the cutoff.
@@
$K_{ij}=\int x_{i}x_{j}\mathrm{e}^{-V(x)}\mathrm{d}\mu(x)-|m|^{2}\quad\text{where}\quad m_{i}=\int x_{i}\mathrm{e}^{-V(x)}\mathrm{d}x.$
The best available bound is logarithmic : $\lambda_{1}^{-1}\asymp\sqrt{\log(d)}r(K)$, see Klartag (2023).
@@ 1.4. Positively curved diffusions, curvature product condition, and cutoff
Recall that the recent result (Salez, 2025, Cor. 1) states that for nonnegatively curved diffusions, and for total variation distance, the product condition implies cutoff at the mixing time (a direct corollary of a theorem bounding the mixing time window). By combining the approaches used in the proof of (Salez, 2025, Cor. 1), the proof of our Corollary 1.4, and the proof of (Chen and Saloff-Coste, 2008, Th. 3.3), we arrive at the following theorem. We state it in the restricted Euclidean setting for simplicity, but it holds in the Riemannian setting, since it is the curvature bound and the diffusion property that matter.
@@
Let $(X_{t}^{(d)})$ be a sequence of diffusion processes on $\mathbb{R}^{d}$ of the form (1.1), with convex potential $V^{(d)}$, initial condition $x_{0}^{(d)}$, and invariant measure $\mu^{(d)}$. Let $S^{(d)}\subset\mathbb{R}^{d}$ be an arbitrary non-empty set of initial conditions. Let $\eta\in(0,1)$ be an arbitrary fixed threshold which does not dependent on $d$. Suppose that there exists positive constants $\kappa^{(d)}$ such that $\mathrm{Hess}(V^{(d)})\geq\kappa_{d}I_{d}$ and that the following curvature product condition holds:
$\lim_{d\to\infty}\kappa^{(d)}t_{0}^{(d)}=+\infty\quad\text{where}\quad t_{0}^{(d)}:=\inf\bigr\{t\geq 0:\sup_{x_{0}^{(d)}\in S^{(d)}}\mathrm{d}_{\mathrm{TV}}(X_{t}^{(d)},\mu^{(d)})\leq\eta\bigr\}.$
@@ 1.5. Rigidity and Gaussian factorization for curved diffusions
$\int\bigr(\langle\mathrm{Hess}(V)\nabla,\nabla f\rangle-\rho|\nabla f|^{2}\bigr)\mathrm{d}\mu=\int\bigr(\langle\mathcal{L}\nabla f-\nabla\mathcal{L}f,\nabla f\rangle-\rho|\nabla f|^{2}\bigr)\mathrm{d}\mu,$ (1.30)
@@
$\int\bigr(\underbrace{\langle\mathrm{Hess}(V)\nabla,\nabla f\rangle-\rho|\nabla f|^{2}}_{\geq 0\text{ by (C3)}}\bigr)\mathrm{d}\mu=(\lambda_{1}-\rho)\int|\nabla f|^{2}\mathrm{d}\mu-\int\|\mathrm{Hess}(f)\|_{\mathrm{HS}}^{2}\mathrm{d}\mu,$ (1.32)
which gives (1.29) [^2] We could also invoque the Brascamp – Lieb inequality $\int f^{2}\mathrm{d}\mu-(\int f\mathrm{d}\mu)^{2}\leq\int\langle\mathrm{Hess}(V)^{-1}\nabla f,\nabla f\rangle\mathrm{d}\mu$, or equivalently the covariance representation of Hörmander – Helffer – Sjöstrand.. Equality in (1.29) is obviously achieved for the OU case $V=\frac{\rho}{2}\left|\cdot\right|^{2}$. More generally, beyond the OU case, we say that the process is rigid when the spectral gap matches the curvature lower bound : $\lambda_{1}=\rho$. Rigidity can be reformulated, and it turns out that for all $\rho>0$, the following items are equivalent:
@@
It is immediate to check that (R2) gives (R1). Conversely, if (R1) is satisfied, then for $f$ such that $\mathcal{L}f=\lambda_{1}f$, we obtain that both sides of (1.32) are identically zero. This gives $\mathrm{Hess}(f)\equiv 0$, hence $f$ is affine, and this also gives that the constant vector $\nabla f$ is in the kernel of the symmetric positive semidefinite matrix $\mathrm{Hess}(V)-\rho\mathrm{I}_{d}$. Moreover $\nabla f\neq 0$ since $f$ is not constant, leading to (R2). We observe that by iterating the proceduce, we get a Gaussian factor with same dimension as the eigenspace of $\lambda_{1}$.
@@ 1.6. Riemannian manifolds
When the manifold is unweighted, this is the dimension-free version of the famous spectral gap bound due to Lichnérowicz (1958). Actually the Bakry – Émery approach implies stronger functional inequalities, such as a logarithmic Sobolev inequality, and a Gaussian isoperimetric inequality. Note that following for instance (Gigli et al., 2015, Prop. 6.7), when $\rho>0$ the spectrum of $\mathcal{L}$ is discrete and the spectral gap $\lambda_{1}$ is always an eigenvalue of -$\mathcal{L}$.
@@
It order to formulate a neat abstract analogue of Theorem 1.1, we shall use vector-valued functions whose coordinates are eigenfunctions, as defined in the following:
@@
Let $E_{1}$ be the eigenspace of $\mathcal{L}$ associated with the eigenvalue $\lambda_{1}$, in $L^{2}(\mu)$. Let $k_{1}=\dim{E_{1}}$. We define a multi-eigenfunction as being a map $M\to\mathbb{R}^{k_{1}}$ whose coordinates are orthogonal elements of $E_{1}$ in $L^{2}(\mu)$. We denote by $F_{1}$ the set of all multi-eigenfunctions, and by $\mathbb{S}_{F_{1}}$ the set of multi-eigenfunctions whose coordinates are elements of the unit sphere of $E_{1}$ in $L^{2}(\mu)$, thus orthonormal.
The link between this definition and Theorem 1.1 is that the component $\sum_{i=1}^{d}|f_{i}(x_{0})|^{2}$ in the definition of $\Lambda$ in Theorem 1.1 is exactly the $\ell^{2}$ norm of $T(x_{0})$ where $T$ is some element of $\mathbb{S}_{F_{1}}$. The Riemannian analog of Theorem 1.1 is now the following.
@@
$\frac{\mathrm{e}^{-\rho t}}{\sqrt{\rho}}\Bigr(\sup_{x_{0}\in S}\Lambda(x_{0})\Bigr)^{1/2}\leq\mathrm{W}_{2}(X_{t},\mu)\leq\mathrm{e}^{-\rho t}\sup_{x_{0}\in S}\Bigr(\int{d_{M}(x_{0},x)^{2}\mathrm{d}\mu(x)}\Bigr)^{1/2},$
@@
$\Lambda(x)=k_{1}+\sup_{T\in\mathbb{S}_{F_{1}}}|T(x)|_{M}^{2}.$ (1.42)
Moreover, if $S=\{x\in M:d_{M}(x,m)\leq R\}$, then we can replace $\sup_{x_{0}\in S}\Lambda(x_{0})$ by
$\sup_{T\in\mathbb{S}_{F_{1}}}(|T(m)|_{M}^{2}+\rho R^{2}).$ (1.43)

### v4

# On cutoff via rigidity for high-dimensional curved diffusions

@@ Abstract.
We consider overdamped Langevin diffusions in Euclidean space, with curvature equal to the spectral gap. This includes the Ornstein–Uhlenbeck process as well as non-Gaussian and non-product extensions with convex interaction, such as the Dyson process from random matrix theory. We show that a cutoff phenomenon or abrupt convergence to equilibrium occurs in high dimension, at a critical time equal to the logarithm of the dimension divided by twice the spectral gap. This cutoff holds for Wasserstein distance, total variation, relative entropy, and Fisher information. A key observation is a relation to a spectral rigidity, linked to the presence of a Gaussian factor. A novelty is the extensive usage of functional inequalities, even for short-time regularization, and the reduction to Wasserstein. The proofs are short and conceptual. Since the product condition is satisfied, an Lp cutoff holds for all p. We moreover discuss a natural extension to Riemannian manifolds, a link with logarithmic gradient estimates in short-time for the heat kernel, and ask about stability by perturbation. Finally, beyond rigidity but still for diffusions, a discussion around the recent progress on the product condition for non-negatively curved diffusions leads us to introduce a new curvature product condition.
@@ 1.1. Diffusions on Euclidean spaces, with convex potential
1.2. Cutoff for high-dimensional curved diffusions
@@
$\mathrm{W}_{2}(\nu,\mu)=\inf_{(X,Y)}\sqrt{\mathbb{E}(d(X,Y)^{2})}$ (1.4)
where the infimum runs over all couples $(X,Y)$ of random variables with $X\sim\nu$ and $Y\sim\mu$, see Santambrogio (2015). We also write $\mathrm{W}_{2}(Z,\mu)=\mathrm{W}_{2}(\mathrm{Law}(Z),\mu)$. We take d(x,y)=|xy|d(x,y)=|x-y| on $\mathbb{R}^{k}$.
@@
$\frac{\mathrm{e}^{-2\lambda_{1}t}}{\lambda_{1}}\sup_{x_{0}\in S}\Lambda(x_{0})+\frac{k_{1}\bigl(1-\sqrt{1-\mathrm{e}^{-2\lambda_{1}t}}\bigr)^{2}}{\lambda_{1}}\leq\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t},\mu)^{2}\leq\mathrm{e}^{-2\lambda_{1}t}\sup_{x_{0}\in S}\int|x-x_{0}|^{2}\mathrm{d}\mu(x)$ (1.6)
with, denoting $E_{1}$ the eigenspace of -$\mathcal{L}$ associated to $\lambda_{1}$ and $k_{1}=\dim(E_{1})$,
$\Lambda(x_{0})=\sup_{(f_{1},\ldots,f_{k_{1}})}\sum_{i=1}^{k_{1}}|f_{i}(x_{0})|^{2},$ (1.7)
where the supremum runs over the set of orthonormal bases of $E_{1}$.
Moreover, if $\mu$ is centered and $S=\{x\in\mathbb{R}^{d}:|x|\leq R\}$, then $\sup_{x_{0}\in S}\Lambda(x_{0})$ can be replaced by $\lambda_{1}R^{2}$.
Remark 1.2 ($k_{1}$ and $\Lambda(x_{0})$).
We always have $k_{1}\leq d$, see (Barthe and Klartag, 2020, Lem. 14). The factor involving $k_{1}$ could be dropped in the lower bound, since it is nonnegative. This will be good enough for the proof of Corollary 1.3 below, but there are situations, such as very small sets of initial data, where the extra factor is useful to get sharp bounds.
By elliptic regularity, the elements of $E_{1}$ are $\mathcal{C}^{2}$ and in particular continuous, see (Evans, 2010, Ch. 6). Moreover, since $E_{1}$ is finite dimensional, the linear map $f\in E_{1}\mapsto\ell(f)=f(x_{0})$ is continuous, and by the Riesz representation theorem, there exists an element $h_{x_{0}}\in E_{1}$ such that $\ell=\langle\cdot,h_{x_{0}}\rangle$, and then the Parseval identity gives $\sum_{i=1}^{k_{1}}|f_{i}(x_{0})|^{2}=\sum_{i=1}^{k_{1}}|\langle f_{i},h_{x_{0}}\rangle|^{2}=\|h_{x_{0}}\|^{2}$ for all orthonormal basis $f_{1},\ldots,f_{k_{1}}$ of $E_{1}$. Therefore the supremum in the definition of $\Lambda(x_{0})$ is useless.
@@
$S^{(d)}=\{x\in\mathbb{R}^{d}:|x-m^{(d)}|\leq c\sqrt{d}\},\quad\text{for an arbitrary constant $c>0$}.$ (1.11)
The assumption $c>0$ is essential to get a cutoff at critical time $\log(d)/(2\lambda_{1}^{(d)})$. Indeed, in the case of the $d$-dimensional Ornstein–Uhlenbeck process with $S^{(d)}=\{m^{(d)}\}$, one has a cutoff at critical time $\log(d)/(4\lambda_{1}^{(d)})$, and this is an effect of the second part of the left hand side of (1.6).
@@
$\mathrm{H}(\nu\mid\mu)=\int\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\log\Bigl(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\Bigr)\mathrm{d}\mu=\int\log\Bigl(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\Bigr)\mathrm{d}\nu\in[0,+\infty]$ (1.13)
@@
$\mathrm{I}(\nu\mid\mu)=\int\frac{|\nabla\frac{\mathrm{d}\nu}{\mathrm{d}\mu}|^{2}}{\frac{\mathrm{d}\nu}{\mathrm{d}\mu}}\mathrm{d}\mu=\int\Bigl|\nabla\log\Bigl(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\Bigr)\Bigr|^{2}\mathrm{d}\nu\in[0,+\infty],$ (1.14)
with the convention $\mathrm{I}(\nu\mid\mu)=+\infty$ if $\nu$ is not absolutely continuous with respect to $\mu$. By an abuse of notation, we take the freedom of replacing $\nu$ by $X\sim\nu$ in these expressions.
@@
Under the setting of Corollary 1.3, and if additionally
$\lambda_{1}^{(d)}=o_{d\to\infty}(\log(d))$ (1.15)
then, for all $\varepsilon>0$,
@@
If $\pi$ and $\pi^{\perp}$ are the orthogonal projections on $\mathbb{R}(1,\ldots,1)$ and its orthogonal, respectively, then $|x|^{2}=|\pi(x)|^{2}+|\pi^{\perp}(x)|^{2}$, while the translation invariance of $W$ in the direction $(1/\sqrt{d},\ldots,1/\sqrt{d})$ gives $W(x)=W(\pi(x)+\pi^{\perp}(x))=W(\pi^{\perp}(x))$, therefore
@@
This approach can be adapted to the special degenerate case of the Dyson–OU or DOU process studied in Chafaï and Lehec (2020); Boursier et al. (2023), when
@@
the degeneracy being equivalent to define the Dyson–OU process on the convex domain $\{x\in\mathbb{R}^{d}:x_{1}>\cdots>x_{d}\}$ instead of on the whole space $\mathbb{R}^{d}$, to exploit convexity. In this case, the symmetric Hermite polynomial $x_{1}+\cdots+x_{d}$ is an eigenfunction associated to the spectral gap, the tip of an iceberg of integrability, as observed in Lassalle (1991).
@@ 1.3. Product condition and $L^{p}$ cutoff at the mixing time
$\lim_{d\to\infty}\sup_{x^{(d)}_{0}\in S^{(d)}}\Bigl\|\frac{\mathrm{d}\mu^{(d)}_{t_{d}}}{\mathrm{d}\mu^{(d)}}-1\Bigr\|_{\mathrm{L}^{p}(\mu^{(d)})}=\begin{cases}\max&\text{if $t_{d}=(1-\varepsilon)t_{*}$}\\ 0&\text{if $t_{d}=(1+\varepsilon)t_{*}$}\end{cases}$ (1.25)
where $\max=2$ if $p=1$ and $\max=+\infty$ if $p>1$, see Chafaï (2024) for a pedagogical presentation. We emphasize that Corollaries 1.31.4, and 1.6 give the mixing time as well as the cutoff.
@@
$K_{ij}=\int x_{i}x_{j}\mathrm{d}\mu(x)-m_{i}m_{j}\quad\text{where}\quad m_{i}=\int x_{i}\mathrm{d}\mu(x).$
The best available upper bound is $\lambda_{1}^{-1}\leq c_{d}r(K)$ with $c_{d}$ logarithmic in $d$, see Klartag (2023). The reverse bound $r(K)\leq\lambda_{1}^{-1}$ is always available.
@@ 1.4. Positively curved diffusions, curvature product condition, and cutoff
Recall that the recent result (Salez, 2025, Cor. 1) states that for non-negatively curved diffusions, and for total variation distance, the product condition implies cutoff at the mixing time (a direct corollary of a theorem bounding the mixing time window). By combining the approaches used in the proof of (Salez, 2025, Cor. 1), the proof of our Corollary 1.4, and the proof of (Chen and Saloff-Coste, 2008, Th. 3.3), we arrive at the following theorem. We state it in the restricted Euclidean setting for simplicity, but it holds in the Riemannian setting, since it is the curvature bound and the diffusion property that matter.
@@
Let $(X_{t}^{(d)})$ be a sequence of diffusion processes on $\mathbb{R}^{d}$ of the form (1.1), with convex potential $V^{(d)}$, initial condition $x_{0}^{(d)}$, and invariant measure $\mu^{(d)}$. Let $S^{(d)}\subset\mathbb{R}^{d}$ be an arbitrary non-empty set of initial conditions. Let $\eta\in(0,1)$ be an arbitrary fixed threshold which does not depend on $d$. Suppose that there exists positive constants $\kappa^{(d)}$ such that $\mathrm{Hess}(V^{(d)})\geq\kappa^{(d)}I_{d}$ and that the following curvature product condition holds:
$\lim_{d\to\infty}\kappa^{(d)}t_{0}^{(d)}=+\infty\quad\text{where}\quad t_{0}^{(d)}:=\inf\bigl\{t\geq 0:\sup_{x_{0}^{(d)}\in S^{(d)}}\mathrm{d}_{\mathrm{TV}}(X_{t}^{(d)},\mu^{(d)})\leq\eta\bigr\}.$
@@ 1.5. Rigidity and Gaussian factorization for curved diffusions
$\int\bigl(\langle\mathrm{Hess}(V)\nabla f,\nabla f\rangle-\rho|\nabla f|^{2}\bigr)\mathrm{d}\mu=\int\bigl(\langle\mathcal{L}\nabla f-\nabla\mathcal{L}f,\nabla f\rangle-\rho|\nabla f|^{2}\bigr)\mathrm{d}\mu,$ (1.31)
@@
$\int\bigl(\underbrace{\langle\mathrm{Hess}(V)\nabla f,\nabla f\rangle-\rho|\nabla f|^{2}}_{\geq 0\text{ by (C3)}}\bigr)\mathrm{d}\mu=(\lambda_{1}-\rho)\int|\nabla f|^{2}\mathrm{d}\mu-\int\|\mathrm{Hess}(f)\|_{\mathrm{HS}}^{2}\mathrm{d}\mu,$ (1.33)
which gives (1.30) [^2] We could also invoke the Brascamp–Lieb inequality $\int f^{2}\mathrm{d}\mu-(\int f\mathrm{d}\mu)^{2}\leq\int\langle\mathrm{Hess}(V)^{-1}\nabla f,\nabla f\rangle\mathrm{d}\mu$, or equivalently the covariance representation of Hörmander–Helffer–Sjöstrand.. Equality in (1.30) is obviously achieved for the OU case $V=\frac{\rho}{2}\left|\cdot\right|^{2}$. More generally, beyond the OU case, we say that the process is rigid when the spectral gap matches the curvature lower bound: $\lambda_{1}=\rho$. Rigidity can be reformulated, and it turns out that for all $\rho>0$, the following items are equivalent:
@@
It is immediate to check that (R2) gives (R1). Conversely, if (R1) is satisfied, then for $f$ such that $\mathcal{L}f=\lambda_{1}f$, we obtain that both sides of (1.33) are identically zero. This gives $\mathrm{Hess}(f)\equiv 0$, hence $f$ is affine, and this also gives that the constant vector $\nabla f$ is in the kernel of the symmetric positive semidefinite matrix $\mathrm{Hess}(V)-\rho\mathrm{I}_{d}$. Moreover $\nabla f\neq 0$ since $f$ is not constant, leading to (R2). We observe that by iterating the procedure, we get a Gaussian factor with same dimension as the eigenspace of $\lambda_{1}$.
@@ 1.6. Riemannian manifolds
When the manifold is unweighted, this is the dimension-free version of the famous spectral gap bound due to Lichnérowicz (1958). Actually the Bakry–Émery approach implies stronger functional inequalities, such as a logarithmic Sobolev inequality, and a Gaussian isoperimetric inequality. Note that following for instance (Gigli et al., 2015, Prop. 6.7), when $\rho>0$ the spectrum of $\mathcal{L}$ is discrete and the spectral gap $\lambda_{1}$ is always an eigenvalue of -$\mathcal{L}$. Moreover, the eigenfunctions associated to the spectral gap are continuous.
@@
In order to formulate a neat abstract analogue of Theorem 1.1, we shall use vector-valued functions whose coordinates are eigenfunctions, as defined in the following:
@@
Let $E_{1}$ be the eigenspace of -$\mathcal{L}$ associated with the eigenvalue $\lambda_{1}$, in $L^{2}(\mu)$. Let $k_{1}=\dim{E_{1}}$. We define a multi-eigenfunction as being a map $M\to\mathbb{R}^{k_{1}}$ whose coordinates are orthogonal elements of $E_{1}$ in $L^{2}(\mu)$. We denote by $F_{1}$ the set of all multi-eigenfunctions, and by $\mathbb{S}_{F_{1}}$ the set of multi-eigenfunctions whose coordinates are elements of the unit sphere of $E_{1}$ in $L^{2}(\mu)$, thus orthonormal.
The link between this definition and Theorem 1.1 is that the component $\sum_{i=1}^{k_{1}}|f_{i}(x_{0})|^{2}$ in the definition of $\Lambda$ in Theorem 1.1 is exactly the $\ell^{2}$ norm of $T(x_{0})$ where $T$ is some element of $\mathbb{S}_{F_{1}}$. The Riemannian analog of Theorem 1.1 is now the following.
@@
$\frac{\mathrm{e}^{-2\rho t}}{\rho}\sup_{x_{0}\in S}\Lambda(x_{0})+\frac{k_{1}\bigl(1-\sqrt{1-\mathrm{e}^{-2\rho t}}\bigr)^{2}}{\rho}\leq\sup_{x_{0}\in S}\mathrm{W}_{2}(X_{t},\mu)^{2}\leq\mathrm{e}^{-2\rho t}\sup_{x_{0}\in S}\int d_{M}(x_{0},x)^{2}\mathrm{d}\mu(x),$ (1.43)
@@
$\Lambda(x)=\sup_{T\in\mathbb{S}_{F_{1}}}|T(x)|^{2}.$ (1.44)
Moreover, if $S=\{x\in M:d_{M}(x,m)\leq R\}$ with $\Lambda(m)=0$ then we can replace $\sup_{x_{0}\in S}\Lambda(x_{0})$ by $\rho R^{2}$.